OpenStudy (anonymous):

19 gms of molten SnCl2 is electrolysed for sometime using inert electrodes until 0.119 gms of Sn is deposited at cathode. No substance lost during electrolysis. Find the ratio of Sncl2 : SnCl4

OpenStudy (anonymous):

@mos1635 - can you help me out with this if you can?

OpenStudy (anonymous):

1:2

OpenStudy (anonymous):

noway :) its 71.34 : 1

OpenStudy (anonymous):

but how do you do it? the way i did it was - No. of moles of SnCl2 before reaction = 19/187.5 = 0.1001 moles; No of moles of Sn discharged = 0.119/119 = 0.001 moles; No of moles of Sn remaining in the solution = 0.1001 - 0.001 = 0.099 moles; therefore, same no. of moles of Sncl4 will be produced. therefore mass of Sncl4 produced = atomic mass x no.of moles = 261 x 0.099 = 25.8 gms. therefore ratio = 19/25.8 = 0.73 :1 which is wrong

OpenStudy (mos1635):

i can not get the reaction........ i suppose that prodused Cl oxidize part of SnCl2 to SnCl4.....is that wrigt??/

OpenStudy (anonymous):

Sncl2 ----> Sn + SnCl4

OpenStudy (mos1635):

2 Sncl2 ----> Sn + SnCl4 n1 0 0 -n2 +n2/2 +n2/2 n1-n2 n2/2 n2/2

OpenStudy (mos1635):

please check moleculare masses

OpenStudy (mos1635):

n1=19/187.5 n2/2=0.119/119 =>n2=0.002 asked ratio (n1-n2)*187.5/(n2/2)*261 =71.36:1

OpenStudy (mos1635):

that is an exact resalt with no entermidiate aproximations

OpenStudy (mos1635):

i was thinking about that all night in cases like this where so exact resalts are asked 71.34 : 1 we mast give the exact constants we are suppose to use for example i use for Cl MM=35.5 sor Sn 119 there for my initial MM for SnCl was 120 in your case i have used your given values for MMs as you have posted above.